24. Solve recurrence relation an+1-5+2an for n20 given a01.


1
Expert's answer
2021-07-06T14:45:02-0400

Solve recurrence relation an+1-5+2an for n20 given a01.

Dividing by a suitable function for step 2 to work

an 2 n = an−1 2 n−1 + 1 2 n .

sn = sn−1 + . . . .

sn = sn−1 + 1 2 n where sn = an 2 n .

an = 2n sn = 2n − 1.

Check that an=2n+1an=2n+1 is the solution to the recurrence relation an=2an−1−1an=2an−1−1 with a1=3.

3